# MFTCXplain: A Multilingual Benchmark Dataset for Evaluating the Moral Reasoning of LLMs through Hate Speech Multi-hop Explanations

### v3

@@ 6 Evaluation
For the final outcome, we obtain multi-hop explanations that include the hate label, the moral sentiments, and the rationales for those moral sentiments, all predicted by the LLMs. Table 7 presents the evaluation results. We evaluated the performance of hate speech classification using the standard F1 score. For moral sentiment prediction, however, we employed an adapted F1 score, which considers an LLM’s prediction correct if it matches any of the one, two, or three moral sentiment labels assigned by human annotators for the same tweet. In addition, to assess whether LLMs can provide human-aligned rationales for moral classifications, we used plausibility and linguistic quality, which are state-of-the-art metrics to assess how well the rationales provided by models align with human rationales DeYoung et al. (2020). Specifically, for plausibility, we used the Jaccard lexical overlap metric (Jaccard, 1901), and for linguistic quality, we used the BERTScore semantic similarity metric (Zhang et al., 2020). We also introduce a customized version of these metrics, as our dataset includes human-annotated rationales for multiple MFT categories, as detailed below.
@@
$\text{Jaccard}_{\text{bid}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{2}\left(\text{Jaccard}(\text{Pred}_{i}\rightarrow\text{Gold}_{i})+\text{Jaccard}(\text{Gold}_{i}\rightarrow\text{Pred}_{i})\right)$ (2)
@@
$\text{BERTScore}_{\text{max}}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}\,\text{BERTScore}(\text{LLM}_{i},\text{Human}_{ij})$ (3)
@@
$\text{BERTScore}_{\text{bid}}=\frac{1}{2}\left(\text{AvgMax}_{\text{LLM}\rightarrow\text{Human}}+\text{AvgMax}_{\text{Human}\rightarrow\text{LLM}}\right)$ (4)
7 Results and Discussion
As shown in Appendix D Figure 8, when compared to the avg length of human-annotated rationales, we observe that model-generated rationales are substantially more verbose. Notably, GPT models are less verbose overall than LLaMA models.
@@ 7 Results and Discussion
LLaMA-70B zero-shot (EN) 0.739 0.324 0.347 0.401 0.431 0.372 0.426 0.578
LLaMA-70B zero-shot (FA) 0.695 0.326 0.313 0.378 0.352 0.594 0.599 0.704
LLaMA-70B zero-shot (PT) 0.642 0.264 0.287 0.481 0.557 0.354 0.438 0.628
LLaMA-70B zero-shot (IT) 0.836 0.340 0.367 0.436 0.438 0.397 0.393 0.532
LLaMA-70B four-shot (EN) 0.636 0.301 0.147 0.352 0.395 0.259 0.338 0.459
LLaMA-70B four-shot (FA) 0.658 0.335 0.231 0.260 0.232 0.553 0.599 0.648
LLaMA-70B four-shot (PT) 0.751 0.264 0.339 0.490 0.414 0.361 0.384 0.489
LLaMA-70B four-shot (IT) 0.787 0.378 0.254 0.333 0.300 0.401 0.414 0.466
@@
GPT-4o zero-shot (EN) 0.738 0.254 0.300 0.419 0.474 0.400 0.451 0.596
GPT-4o zero-shot (FA) 0.642 0.231 0.266 0.363 0.361 0.590 0.599 0.687
GPT-4o zero-shot (PT) 0.624 0.265 0.235 0.428 0.496 0.279 0.369 0.552
GPT-4o zero-shot (IT) 0.826 0.384 0.414 0.478 0.499 0.457 0.463 0.571
GPT-4o four-shot (EN) 0.635 0.225 0.253 0.366 0.419 0.301 0.375 0.496
GPT-4o four-shot (FA) 0.744 0.275 0.261 0.389 0.400 0.601 0.611 0.684
GPT-4o four-shot (PT) 0.760 0.317 0.243 0.305 0.262 0.249 0.226 0.293
GPT-4o four-shot (IT) 0.680 0.390 0.294 0.384 0.346 0.408 0.413 0.479
@@
Furthermore, as shown in Table 7 and Figures 3 and 4, LLMs demonstrate considerably higher performance on binary hate speech classification compared to the task of moral sentiment prediction. F1 scores for hate speech detection frequently exceed 0.7, reaching up to 0.836 for LLaMA-70B (IT), while moral sentiment prediction consistently lags behind, rarely surpassing 0.35. This disparity underscores the complexity of moral reasoning tasks, which require not just surface-level understanding but contextual and culturally situated interpretation. Notably, Italian appears to be the most accessible language for both models, especially for GPT-4o, which achieves its highest moral F1 (0.440) in this language under CoT prompting. In contrast, Persian and Portuguese consistently show lower performance, reflecting the challenges that even state-of-the-art LLMs face when reasoning over underrepresented languages [^5] The particularly low performance for Persian may also stem from challenges associated with its non-Latin script, which has been shown to limit model alignment in previous work (Delbari et al., 2024a)..
When evaluating the alignment between model-generated rationales and human annotations, performance varies widely across languages and prompting strategies. Despite moderate lexical overlaps in Jaccard metrics (e.g., $\text{Jaccard}_{\text{maxpair}}$ reaching up to 0.565 for LLaMA-70B CoT on PT), semantic similarity metrics such as BERTScore suggest a deeper gap in the models’ capacity to reproduce human moral justifications. For example, GPT-4o under CoT prompting achieves a Jaccard of only 0.099 and BERTScore of 0.198 for Portuguese, revealing a severe misalignment in token-level rationales. Interestingly, while few-shot prompting was expected to improve alignment, it did not consistently outperform zero-shot or CoT prompting across tasks, particularly for rationale generation. These findings highlight the limitations of current LLMs in generating culturally-aware moral explanations and point to the need for high-quality, multilingual moral annotated data as the proposed MFTCXplain.
Finally, according to the cross-cultural corpus analysis, hate speech comprises moralized content. Across English, Italian, Portuguese, and Persian, tweets labeled as hate speech rarely appeared as “Non-Moral” and frequently relied on moral foundations (e.g., Harm, Degradation, Fairness, and Loyalty. This supports prior findings that hate speech is framed by morality (Kennedy et al., 2023).
@@ 8 Conclusion
This paper introduced MFTCXplain, the first multilingual benchmark designed to evaluate the moral reasoning of large language models through structured, multi-hop hate speech explanations grounded in Moral Foundations Theory. By incorporating expert-annotated rationales across four linguistically and culturally diverse languages, our dataset enables a fine-grained and interpretable analysis of moral reasoning in LLMs. Empirical findings reveal significant misalignments between LLM-generated justifications and human moral reasoning, underscoring the limitations of current LLM reasoning techniques. Our cross-cultural corpus analysis also reveals that hate speech is deeply moralized and is rarely used without moral sentiment. MFTCXplain thus provides not only a new approach for assessing model behavior, but also a critical resource for driving progress toward more transparent, culturally aware, and ethically aligned NLP systems. We hope our dataset and findings can contribute to ongoing and future research in evaluating the multilingual moral reasoning of LLMs, hate speech detection, and explainability in natural language processing and machine learning.
@@
Acknowledgements
This work was partially supported by a grant from Amazon. The authors thank Dr. Morteza Dehghani and the Morality and Language Lab at USC for their early guidance and support on this project.
@@ Appendix A Appendix
A.1 Moral Foundations Coding Guide (Updated) [^6] This guide is based on the the original Moral Foundations Coding Guide which can be found in the Appendix of Hoover et al. (2020b).
Moral expressions in text serve as informationally rich indicators of individuals’ moral values. Whether individuals are signaling their moral beliefs or concerns, framing particular issues or events in moral terms, or expressing a moral emotion, moral expressions are a domain of human language which can inform as to the nature of morality (Atari et al., 2023a). Here, we describe a taxonomy and set of instructions for annotating moral content in natural language, based on Moral Foundations Theory. This taxonomy can be used for the annotation of individual Tweets, Facebook posts, other social media, transcribed speech, and other textual media.
In this coding guide, we describe the theoretical framework that we rely on to operationalize moral values, Moral Foundations Theory (MFT; Haidt and Graham, 2007; Graham et al., 2013), describe how moral expressions are annotated, and provide detailed examples and procedures for the process of annotation.
@@ A.1.2 Instructions for Annotators: Annotating moral concerns in language
Table 8: Five moral foundations, which map to five domains of moral language
Foundation Description Example Item
Care Intuitions about avoiding emotional and physical damage to another individual. It underlies virtues of kindness, gentleness, and nurturing. I believe that compassion for those who are suffering is one of the most crucial virtues.
Fairness Intuitions about egalitarian treatment and equal outcome for all individuals and groups. It underlies virtues of social justice and equality. The world would be a better place if everyone made the same amount of money
Loyalty Intuitions about cooperating with ingroups and competing with outgroups. It underlies virtues of patriotism and self-sacrifice for the group. I believe the strength of a sports team comes from the loyalty of its members to each other
Authority Intuitions about deference toward legitimate authorities and high-status individuals. It underlies virtues of leadership and respect for tradition. I think obedience to parents is an important virtue
Purity Intuitions about avoiding bodily and spiritual contamination and degradation. It underlies virtues of sanctity, nobility, and cleanliness. It underlies the widespread idea that the body is a temple that can be desecrated by immoral activities and contaminants (an idea not unique to religious traditions)
@@ A.2 Annotator Metadata
A total of five annotators contributed to the annotation of this corpus, including three females and two males, all of whom were graduate students (Master’s or PhD level) from participating universities, with a mean age of 29.
For the Portuguese dataset, we had two annotators drawn from distinct cultural zones of Brazil (São Paulo and Minas Gerais), representing different racial groups (Black and White, categorized as “color” in Brazil). For the other languages (Persian, Italian, English), one annotator per language was involved.
@@ A.3 Dataset Sources
For the Portuguese data in our corpus, we collected data from Twitter during the public resignation of the Minister of Justice under the Bolsonaro government on April 24, 2020  [^8] https://www.cnnbrasil.com.br/politica/sergio-moro-pede-demissao-do-governo-bolsonaro/. We collected 6,000 tweets precisely during the public resignation announcement. For the MFTCxplain, we randomly selected 542 tweets with hate speech or offensive language and 509 tweets with non-hate speech or offensive language. The hate speech annotation was performed by a linguist who is also an expert in hate speech, based on the annotation schema provided in Vargas et al. (2022)
@@ B.1 Linguistic Analysis
Figure 6: Top-5 lemmas per moral category across four languages in terms of specificity measure (TXM tool) extracted from posts classified according to human annotation and LLaMA-70B and GPT-4o zero-shot
@@ C.1 Moral Foundations of Hate Speech Across Cultures
To better understand the moral framework of hate speech, we analyze the prevalence of specific moral foundations across English, Italian, Portuguese, and Persian, revealing both shared patterns and culturally distinct narratives, as shown in Figure 7.
@@
Figure 7: Moral Foundation Theory Categories by Hate Speech and Non-Hate Speech Across Languages.
@@
Note that hate speech is a complex, subjective, and context-dependent phenomenon that poses significant theoretical and empirical challenges across cultures (Erjavec and Kovačič, 2012). Legally, the classification of hate speech remains contentious, with diverging frameworks across nations; for example, the United States protects most hate speech under the First Amendment, whereas Germany enforces stringent prohibitions against certain forms of hate speech under its Volksverhetzung laws (Kennedy et al., 2023). These differences create ongoing tensions for content moderation across global platforms, as platforms must balance local legal norms with transnational standards (Fortuna et al., 2020). To better understand the murky linguistic and moral boundaries of hate speech, recent work has argued that hate speech frequently carries an informative underlying tone of moral rhetoric (Kennedy et al., 2023). However, morality itself varies across cultures (Atari et al., 2023a), suggesting that the moral underpinnings of hate speech may also differ cross-culturally.
In our cross-cultural corpus, we observe patterns consistent with previous research on the moral framing of hate (Kennedy et al., 2023). Across English, Italian, Portuguese, and Persian corpora, the moral foundations of Harm and Degradation emerge as the most frequent moral dimensions co-occurring with hate speech, aligning with findings from Nazi propaganda and neo-Nazi discourse analyzed by Kennedy et al. (2023). This moral pattern of hate speech, distinctively invoking moralized narratives of harm, corruption, and danger, remains remarkably consistent, with some variation across languages. For example, Portuguese and Persian tweets often framed hate through violations of fairness and loyalty in addition to harm, whereas Italian and English data more strongly emphasized degradation and betrayal. Notably, in all four languages, tweets labeled as containing hate speech were rarely classified as “Non-Moral,” reinforcing the idea that hate speech across cultures almost always carries some kind of moral framing.
A striking pattern in our analysis is the disproportionate association between hate speech and vice labels rather than virtue labels within each moral foundation. For example, hate speech tweets were much more likely to be tagged with “Harm” (vice) than “Care” (virtue), or with “Degradation” rather than “Purity.” This finding suggests that hate speech is commonly framed around moral violations – depictions of groups as threats, betrayers, or corrupters. However, not all hate speech relies solely on vice framing. In some cases, annotators identified virtue-based moral justifications — for example, tweets expressing loyalty or fairness concerns as a rationale for outgroup hostility (supporting findings from Kennedy et al. (2023)). Additionally, we observed that vice language also appeared in non-hate tweets, reflecting that moral violations can be discussed critically without necessarily expressing hatred toward the individuals or groups involved.
While the overall pattern of vice framing in hate speech was robust, notable differences emerged across languages. In the Persian and English corpora, certain vice categories, particularly Cheating, had more non-hate tweets than hate tweets. In contrast, in the Portuguese and Italian corpora, each vice category was more frequently associated with hate speech than with non-hate speech. This divergence may partly reflect the dominance of Harm-related foundations within the hate-labeled tweets in the English and Persian datasets, where Harm narratives may have overshadowed other vice categories such as Subversion and Cheating. These patterns highlight the importance of considering cultural and contextual factors when interpreting the moral tone of hate across languages.
Moreover, through our annotator-in-the-loop sessions, annotators highlighted the difficulty in distinguishing between vice and virtue framing within the same foundation, particularly in morally charged, culturally nuanced tweets. This ambiguity underscores that hate speech is not only a matter of what is said but how moral narratives are strategically deployed.
Overall, our findings demonstrate that analyzing the moral tone of hate – across diverse cultural contexts – offers crucial insights into how hate speech is constructed, justified, and perpetuated in online discourse.

### v4

@@ 5 Evaluation
For the final outcome, we obtain multi-hop explanations that include the hate label, the moral sentiments, and the rationales for those moral sentiments, all predicted by the LLMs [^6] An ablation study removing multi-hop reasoning and theoretical scaffolding (see the Appendix E) sharply degrades moral classification and rationale quality, supporting the role of structured reasoning.. Table 7 presents the evaluation results [^7] Further analysis of model errors, including per-label false-positive/false-negative counts, is provided in Appendix F.1. We evaluated the performance of hate speech classification using the standard F1 score. For moral sentiment prediction, however, we employed an adapted F1 score, which considers an LLM’s prediction correct if it matches any of the one, two, or three moral sentiment labels assigned by human annotators for the same tweet. In addition, to assess whether LLMs can provide human-aligned rationales for moral classifications, we used plausibility and linguistic quality, state-of-the-art metrics to assess how well the rationales provided by models align with human rationales DeYoung et al. (2020). Specifically, for plausibility, we used the Jaccard lexical overlap metric (Jaccard, 1901), and for linguistic quality, we used the BERTScore semantic similarity metric (Zhang et al., 2020) [^8] Human–human BERTScore ceilings for our task are reported in Appendix H.1; LLM scores should be interpreted relative to this ceiling. We also introduce a customized version of these metrics, as our dataset includes human-annotated rationales for multiple MFT categories, as detailed below.
@@
$\begin{split}\text{Jaccard}_{\text{bid}}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{2}\left(\text{Jaccard}(\text{Pred}_{i}\rightarrow\text{Gold}_{i})+\text{Jaccard}(\text{Gold}_{i}\rightarrow\text{Pred}_{i})\right)\end{split}$ (2)
@@
$\begin{split}\text{BERTScore}_{\text{max}}=\frac{1}{N}\sum_{i=1}^{N}\max_{j}\,\text{BERTScore}(\text{LLM}_{i},\text{Human}_{ij})\end{split}$ (3)
@@
$\begin{split}\text{BERTScore}_{\text{bid}}=\frac{1}{2}\left(\text{AvgMax}_{\text{LLM}\text{Human}}+\text{AvgMax}_{\text{Human}\text{LLM}}\right)\end{split}$ (4)
@@
LLaMA-70B 0-shot (EN) 0.739 0.324 0.347 0.401 0.431 0.372 0.426 0.578
LLaMA-70B 0-shot (FA) 0.695 0.326 0.313 0.378 0.352 0.594 0.599 0.704
LLaMA-70B 0-shot (PT) 0.642 0.264 0.287 0.481 0.557 0.354 0.438 0.628
LLaMA-70B 0-shot (IT) 0.836 0.340 0.367 0.436 0.438 0.397 0.393 0.532
LLaMA-70B 4-shot (EN) 0.636 0.301 0.147 0.352 0.395 0.259 0.338 0.459
LLaMA-70B 4-shot (FA) 0.658 0.335 0.231 0.260 0.232 0.553 0.599 0.648
LLaMA-70B 4-shot (PT) 0.751 0.264 0.339 0.490 0.414 0.361 0.384 0.489
LLaMA-70B 4-shot (IT) 0.787 0.378 0.254 0.333 0.300 0.401 0.414 0.466
@@
GPT-4o 0-shot (EN) 0.738 0.254 0.300 0.419 0.474 0.400 0.451 0.596
GPT-4o 0-shot (FA) 0.642 0.231 0.266 0.363 0.361 0.590 0.599 0.687
GPT-4o 0-shot (PT) 0.624 0.265 0.235 0.428 0.496 0.279 0.369 0.552
GPT-4o 0-shot (IT) 0.826 0.384 0.414 0.478 0.499 0.457 0.463 0.571
GPT-4o 4-shot (EN) 0.635 0.225 0.253 0.366 0.419 0.301 0.375 0.496
GPT-4o 4-shot (FA) 0.744 0.275 0.261 0.389 0.400 0.601 0.611 0.684
GPT-4o 4-shot (PT) 0.760 0.317 0.243 0.305 0.262 0.249 0.226 0.293
GPT-4o 4-shot (IT) 0.680 0.390 0.294 0.384 0.346 0.408 0.413 0.479
@@
6 Results and Discussion
Results are shown in Table 7, and Figures 8, 3, and 4. We find that model-generated rationales are substantially more verbose than human annotations, with GPT models generally more concise than LLaMA models, as shown in Figure 8 in Appendix D).
@@ 6 Results and Discussion
Furthermore, as shown in Table 7 and Figures 3 and 4, LLMs demonstrate considerably higher performance on binary hate speech classification compared to the task of predicting moral sentiment. For example, F1 scores for hate speech detection frequently exceed 0.7, reaching up to 0.836 for LLaMA-70B (IT), while moral sentiment prediction consistently lags behind, rarely exceeding 0.35. This disparity highlights the complexity of moral reasoning tasks, which require not only a surface-level understanding but also a contextual and culturally situated interpretation.
In particular, Italian appears to be the most accessible language for both models, especially for GPT-4o, which achieves its highest moral F1 (0.440) in this language under CoT prompting. In contrast, Persian and Portuguese consistently show lower performance, reflecting the challenges that even state-of-the-art LLMs face when reasoning over underrepresented languages [^9] The particularly low performance for Persian may also stem from challenges associated with its non-Latin script, which has been shown to limit model alignment in previous work (Delbari et al., 2024a)..
When evaluating the alignment between model-generated rationales and human annotations, performance varies widely across languages and prompting strategies. Despite moderate lexical overlaps in Jaccard metrics (e.g., $\text{Jaccard}_{\text{maxpair}}$ reaching up to 0.565 for LLaMA-70B CoT on PT), semantic similarity metrics such as BERTScore suggest a deeper gap in the models’ capacity to reproduce human moral justifications. For example, GPT-4o under CoT prompting achieves a Jaccard of only 0.099 and BERTScore of 0.198 for Portuguese, revealing a severe misalignment in token-level rationales.
Interestingly, while few-shot prompting was expected to improve alignment, it did not consistently outperform zero-shot or CoT prompting across tasks, particularly for rationale generation. These findings highlight the limitations of current LLMs in generating culturally-aware moral explanations and point to the need for high-quality, multilingual moral annotated data as the proposed MFTCXplain [^10] Qualitative human validation of LLM rationales and error modes appears in Appendix G..
Finally, according to the cross-cultural corpus analysis, hate speech consistently comprises moralized content. Across English, Italian, Portuguese, and Persian, tweets labeled as hate speech rarely appeared as “Non-Moral” and frequently relied on moral foundations such as Harm, Degradation, Fairness, and Loyalty. This supports prior findings that hate speech is often framed by moral sentiments (Kennedy et al., 2023).
@@ 7 Conclusion
This paper introduces MFTCXplain, the first multilingual benchmark designed to evaluate the moral reasoning of large language models through structured, multi-hop explanations of hate speech grounded in Moral Foundations Theory. By incorporating expert-annotated rationales across four linguistically and culturally diverse languages, our dataset enables a fine-grained and interpretable analysis of moral reasoning in LLMs. Empirical findings reveal significant misalignments between LLM-generated justifications and human moral reasoning, underscoring the limitations of current LLM reasoning techniques. Our cross-cultural corpus analysis also reveals that hate speech is deeply moralized and is rarely used without moral sentiment. MFTCXplain thus provides not only a new approach for assessing model behavior, but also a critical resource for driving progress toward more transparent, culturally aware, and ethically aligned NLP systems. We hope that our dataset and findings can contribute to ongoing and future research in evaluating the multilingual moral reasoning of LLMs, hate speech detection, and the explainability and interpretability of natural language processing and machine learning.
Acknowledgements
The authors are grateful to São Paulo Research Foundation – FAPESP (grants #2025/01118-2 and #2024/04890-5) for financial support, and Mikel's grant from Amazon. The authors thank Dr. Morteza Dehghani and the Morality and Language Lab at the University of Southern California (USC) for their early guidance and support on this project.
@@ Appendix A Appendix
A.1 Moral Foundations Coding Guide (Updated) [^11] This guide is based on the the original Moral Foundations Coding Guide which can be found in the Appendix of Hoover et al. (2020).
Moral expressions in text serve as informationally rich indicators of individuals’ moral values. Whether individuals are signaling their moral beliefs or concerns, framing particular issues or events in moral terms, or expressing a moral emotion, moral expressions are a domain of human language which can inform as to the nature of morality (Atari et al., 2023a). Here, we describe a taxonomy and set of instructions for annotating moral content in natural language, based on Moral Foundations Theory. This taxonomy can be used for the annotation of individual Tweets, Facebook posts, other social media, transcribed speech, and other textual media. In this coding guide, we describe the theoretical framework that we rely on to operationalize moral values, Moral Foundations Theory (MFT; Haidt and Graham, 2007; Graham et al., 2013), describe how moral expressions are annotated, and provide detailed examples and procedures for the process of annotation.
@@ A.1.2 Instructions for Annotators: Annotating moral concerns in language
Foundation Description Examples
Care Intuitions about avoiding emotional and physical harm to others; underlies virtues of kindness, gentleness, and nurturing. I believe that compassion for those who are suffering is one of the most crucial virtues.
Fairness Intuitions about egalitarian treatment and equal outcomes for all; underlies virtues of social justice and equality. The world would be a better place if everyone made the same amount of money.
Loyalty Intuitions about cooperating with ingroups and competing with outgroups; underlies virtues of patriotism and self-sacrifice for the group. I believe the strength of a sports team comes from the loyalty of its members to each other.
Authority Intuitions about deference toward legitimate authorities and high-status individuals; underlies virtues of leadership and respect for tradition. I think obedience to parents is an important virtue.
Purity Intuitions about avoiding bodily and spiritual contamination or degradation; underlies virtues of sanctity, nobility, and cleanliness. The body is a temple that can be desecrated by immoral activities and contaminants (an idea not unique to religious traditions).
Table 8: Five moral foundations and illustrative example items.
@@
Annotator-in-the-loop details.
For languages with a single primary annotator, we used annotator-in-the-loop procedures emphasizing deliberative discussion, iterative guideline refinement, and spot checks by the research team. Sessions focused on clarifying category boundaries, re-reading borderline cases, and documenting decisions that updated the coding guide. This qualitative pathway provides validity checks in the absence of multiple independent raters.
@@ A.2 Annotator Metadata
A total of five annotators contributed to the annotation of this corpus, including three females and two males, all of whom were graduate students (Master’s or PhD level) from participating universities, with a mean age of 29. For the Portuguese dataset, we had two annotators drawn from distinct cultural zones of Brazil (São Paulo and Minas Gerais), representing different racial groups (Black and White, categorized as “color” in Brazil). For the other languages (Persian, Italian, English), one annotator per language was involved.
@@ A.3 Dataset Sources
For the Portuguese data in our corpus, we collected data from Twitter during the public resignation of the Minister of Justice under the Bolsonaro government on April 24, 2020  [^13] https://www.cnnbrasil.com.br/politica/sergio-moro-pede-demissao-do-governo-bolsonaro/. We collected 6,000 tweets precisely during the public resignation announcement. For the MFTCxplain, we randomly selected 542 tweets with hate speech or offensive language and 509 tweets with non-hate speech or offensive language. The hate speech annotation was performed by a linguist who is also an expert in hate speech, based on the annotation schema provided in Vargas et al. (2022).
@@ B.1 Linguistic Analysis
Figure 6: Top-5 lemmas per moral category across four languages in terms of specificity measure (TXM tool) extracted from posts classified according to human annotation and LLaMA-70B and GPT-4o zero-shot.
@@ C.1 Moral Foundations of Hate Speech Across Cultures
To better understand the moral framework of hate speech, we analyze the prevalence of specific moral foundations in English, Italian, Portuguese, and Persian, revealing both shared patterns and culturally distinct narratives, as shown in Figure 7.
@@
Figure 7: Moral Foundations Theory categories by hate speech and non-hate speech across languages.
@@
Note that hate speech is a complex, subjective and context-dependent phenomenon that poses significant theoretical and empirical challenges across cultures (Erjavec and Kovačič, 2012). Legally, the classification of hate speech remains contentious, with divergent frameworks between nations; for example, the United States protects most hate speech under the First Amendment, while Germany enforces stringent prohibitions against certain forms of hate speech under its Volksverhetzung laws (Kennedy et al., 2023). These differences create ongoing tensions for content moderation, as platforms must balance local legal norms with transnational standards (Fortuna et al., 2020).
To better understand the murky linguistic and moral boundaries of hate speech, recent work has argued that hate speech frequently carries an informative underlying tone of moral rhetoric (Kennedy et al., 2023). However, morality itself varies across cultures (Atari et al., 2023a), suggesting that the moral underpinnings of hate speech may also differ cross-culturally.
In our cross-cultural corpus, we observe patterns consistent with previous research on the moral framing of hate (Kennedy et al., 2023). Across English, Italian, Portuguese, and Persian corpora, the moral foundations of Harm and Degradation emerge as the most frequent moral dimensions co-occurring with hate speech, aligning with findings from Nazi propaganda and neo-Nazi discourse analyzed by Kennedy et al. (2023). This moral pattern of hate speech, distinctively invoking moralized narratives of harm, corruption, and danger, remains remarkably consistent, with some variation across languages. For example, Portuguese and Persian tweets often framed hate through violations of fairness and loyalty in addition to harm, whereas Italian and English data more strongly emphasized degradation and betrayal. Notably, in all four languages, tweets labeled as containing hate speech were rarely classified as “Non-Moral,” reinforcing the idea that hate speech across cultures almost always carries some kind of moral framing. A striking pattern in our analysis is the disproportionate association between hate speech and vice labels rather than virtue labels within each moral foundation. For example, hate speech tweets were much more likely to be tagged with “Harm” (vice) than “Care” (virtue), or with “Degradation” rather than “Purity.” This finding suggests that hate speech is commonly framed around moral violations – depictions of groups as threats, betrayers, or corrupters. However, not all hate speech relies solely on vice framing. In some cases, annotators identified virtue-based moral justifications — for example, tweets expressing loyalty or fairness concerns as a rationale for outgroup hostility (supporting findings from Kennedy et al. (2023)). Additionally, we observed that vice language also appeared in non-hate tweets, reflecting that moral violations can be discussed critically without necessarily expressing hatred toward the individuals or groups involved. While the overall pattern of vice framing in hate speech was robust, notable differences emerged across languages. In the Persian and English corpora, certain vice categories, particularly Cheating, had more non-hate tweets than hate tweets. In contrast, in the Portuguese and Italian corpora, each vice category was more frequently associated with hate speech than with non-hate speech. This divergence may partly reflect the dominance of Harm-related foundations within the hate-labeled tweets in the English and Persian datasets, where Harm narratives may have overshadowed other vice categories such as Subversion and Cheating. These patterns highlight the importance of considering cultural and contextual factors when interpreting the moral tone of hate across languages. Moreover, through our annotator-in-the-loop sessions, annotators highlighted the difficulty in distinguishing between vice and virtue framing within the same foundation, particularly in morally charged, culturally nuanced tweets. This ambiguity underscores that hate speech is not only a matter of what is said but how moral narratives are strategically deployed. Overall, our findings demonstrate that analyzing the moral tone of hate – across diverse cultural contexts – offers crucial insights into how hate speech is constructed, justified, and perpetuated in online discourse.